# A Lensing-Band Approach to Spacetime Constraints

### 11 Dec 2023 v1

@@ II.3 Resulting scheme and numerical implementation
Figure 3: Two schematic examples of $n=1$ lensing bands $\text{LB}^{n=1}(\mathfrak{g})$ (light-shaded regions), for which the respective spacetime $\mathfrak{g}$ remains allowed (left-hand panel) / can be excluded (right-hand panel), after a confident detection of a lensed emission region $\text{LB}^{n\geq 1}_{\text{VLBI}}$ (dark-shaded circular crescent).
From here on, we assume that a lensed emission region (i) has been observed and (ii) must lie, for any given spacetime $\mathfrak{g}$, within its $n=1$ lensing band, $\text{LB}^{n=1}(\mathfrak{g})$. Hence, if the observed lensed emission region $\text{LB}^{n\geq 1}_{\text{VLBI}}$ does not match, or “fit into,” the lensing-band $\text{LB}^{n=1}(\mathfrak{g})$, we can rule out the respective underlying spacetime geometry $\mathfrak{g}$, see Fig. 3 for an schematic representation of this matching procedure. The level of confidence with which the underlying spacetime geometry can be ruled out depends on (i) the level of confidence in the detection of the lensed emission region, and (ii) the level of confidence in the theoretical assumption that it must occur within the lensing band.
We allow for an arbitrary centroid shift by an image-plane vector $(x_{0},y_{0})$ when performing the matching. The centroid shift amounts to the lack of a reference point. Potential observations of the inner shadow [29, 30, 31] or the base of the jet [79] in M87* could provide additional priors for the centroid shift. Examples of such reference points will be shown in Sec. IV.2, but we will not include them in the statistical analysis presented in this work. Such additional priors will only strengthen evidence against the underlying spacetime $\mathfrak{g}$.
In the following, we present an explicit algorithmic scheme with which we decide – given a lensed emission region $\text{LB}^{m\geq n}_{\text{VLBI}}$, as defined in Sec. II.2 – whether or not a given spacetime $\mathfrak{g}$ is excluded or not. The readers not interested in numerical technicalities may skip this section and proceed directly to the results (Sec.III).
@@ II.3.1 Free-floating bisection to obtain the lensing-band region
These conditions evaluate to TRUE whenever the respective image point is to the left of the respective lensing-band boundary $\partial^{\text{(outer)}}\text{LB}^{n}(\mathfrak{g})$ (or $\partial^{\text{(inner)}}\text{LB}^{n}(\mathfrak{g})$), where left and right are defined following the boundary counterclockwise. The only difference between determining the inner and outer boundary is a flip in the condition (or equivalently a flip in clockwise/counterclockwise direction). Other conventions may be chosen and are equivalent to the above.
@@
  • The second major advantage is the adaptive stepsize. The free-floating bisection initializes each subsequent bisection in reference to (at least two) previous points on the boundary. This makes it possible to dynamically adapt the stepsize and precision which, in turn, allows to optimize efficiency while ensuring that the boundary is not lost.
We assume that the free-floating bisection algorithm can be adapted to any other (binary) boundary bisection condition in a 2D plane. Given the generality of this problem and given the above two advantages, the algorithm might be of more general use. We provide further details in Appendix A.
Given this free-floating bisection algorithm, the problem is reduced to standard numerical ray tracing (in curved spacetime $\mathfrak{g}$) at each bisection step. In order to keep the code-development minimal, we use an adapted version of Mathematica’s NDSolve routine for the ray tracing [^5] This is certainly not the most efficient ray tracing and we are inclined to make use of existing optimized ray tracing codes (such as the ones presented in Refs. [81, 15, 82, 83], for example) in future applications.. The ray tracing is performed in Boyer–Lindquist type coordinates. The translation from screen coordinates to ray tracing coordinates is detailed in Appendix C (see also Ref. [80]) and, in fact, only requires that the coordinates converge to oblate spheroidal coordinates at $r\rightarrow\infty$. The observers screen is then placed at $r=10^{5}\,M$ which approximates the $r\rightarrow\infty$ limit sufficiently well [^6] We have tested that the precision of the lensing-band boundary in the Kerr space time at high inclination (see Fig. 5 for some visual examples) remains at least at the subpercent level with the analytical results using the code described in Ref. [61].
@@ II.3.2 Minimizing the uncovered emission region
The normalization ensures that $0\leqslant\mathcal{A}(\mathfrak{g})\leqslant 1$ which is useful for numerical purposes. Once $\mathcal{A}$ is determined, this decides between
@@ II.3.3 Bayesian inference
Given a confident measurement of lensed emission as discussed above, we estimate the parameters of the spacetime $\mathfrak{g}(M,A,p_{1},\ldots,p_{n})$, where $M$ and $A$ denote the asymptotic mass and spin, respectively, and $p_{i}$ denotes extra parameters (e.g., coupling constants or dimensionless “bumpy” parameters [85]), for which the $n=1$ lensing band of an observer at infinity, and at inclination $i$ can cover the considered ring. In other words, we estimate the set of parameters $\vec{\lambda}=\left\{i,M,A,p_{1},\ldots,p_{n}\right\}$ such that its $n=1$ lensing band covers the observed VLBI feature, as Fig. 3 illustrates.
@@
The emcee ensemble of walkers are initialized by sampling from the prior distributions. For all the cases presented in this work, we run until $\sim 10^{5}$ samples are obtained, and burn-in the initial $100$ samples. In general, parameter regions in which the lensing band is wide are easier to find within the posterior likelihood distribution. Recovering the theoretically expected symmetry of all results and correlations under the exchange of $A\leftrightarrow-A$ (as well as $a\leftrightarrow-a$) seems to be another good marker for convergence of the MCMC. The presented results suggest that most (but not all) correlations are fully converged. Given the exploratory character of our work, we do not deem it justified to spend computation time on further convergence.
@@ III Benchmark: Projected constraints in the Kerr spacetime
We do, however, find an interesting correlation, albeit small, between mass and spin: the higher the absolute value of the spin, the larger the inferred mass, cf. $(A,M)$-panel in Fig. 4. The underlying reason for this correlation is a shrinking of the $n=1$ lensing band with increasing spin, cf. left vs right columns in Fig. 5. With very different inference techniques, it has been previously found [87, 39] that such a correlation between mass and spin is more significant in the $n=1$ image than in the critical curve (i.e., in the $n=\infty$ image). It is reassuring that we recover this correlation.
As expected, given that the modelled VLBI feature is a circle, the results disfavour high inclinations. To be specific, the posterior for the projected constraint on inclination results $i\doteqdot{34^{\circ}}^{+28^{\circ}}_{-22^{\circ}}\;$. Indeed, the $n=1$ lensing band becomes very thin as the inclination approaches $i\sim 90^{\circ}$, as shown by the different rows in Fig. 5. We highlight that formally our analysis assumes, and thus all constraints rely on, a confident detection of lensed emission throughout the specified region $\text{LB}^{n\geq 1}_{\text{VLBI}}$, cf. darker and colored regions in Figs. 5 and 3 and the discussion in Section II.2. In particular, the quantitative projected constraints (such as the one on the inclination) which arise from a thinning of the lensing band, implicitly rely on lensed emission being observed not just somewhere but everywhere within the observed region.
@@
Dedicated future studies to determine how confident VLBI observations can constrain the thickness of a lensed emission region are vital to solidify constraints such as the above ones. We also emphasize that the definition of lensing bands (defined with respect to the equatorial plane) is ill-defined at exactly edge-on inclination $i=90^{\circ}$. Respectively, when deforming, e.g., the $n=1$ lensing band with increasing inclination across $i=90^{\circ}$, one encounters a discontinuity, see e.g. [44, Fig. 7]. The quantitative impact of this remains to be scrutinized in future studies and, hence, we are cautious to interpret our results at close to edge-on inclination.
@@ IV Application: Projected constraints on black hole non-uniqueness
  • efficiently explore degeneracies and correlations in large parameter spaces;
  • quantify the constraining power of a confident observation of a lensed emission region;
@@
As in the previous section, we present the results in the form of “projected constraints” on the underlying spacetime parameters.
In Section IV.1, we briefly review parametrized deviations in stationary and axisymmetric spacetimes beyond Kerr, and present a seven-parameter spacetime which we will use as a prototype test case beyond Kerr. In Section IV.2, we then apply the lensing-band framework presented in Section II and discuss the resulting projected constraints.
@@ IV.1 Parametrizing deviations to circular spacetimes
There are various ways in which black-hole candidates may differ from a black hole solution in GR [88]. In recent years, a “theory-agnostic” approach to testing GR has been heavily pursued. This approach involves parametrizing deviations from GR solutions (such as the metric [85] or gravitational potentials [7]) or observables (such as the gravitational waveform [89] or quasinormal modes frequencies [90]).
The exploration of deviations at the level of the metric requires to (i) parametrize a given class of spacetimes in terms of free metric functions, and then (ii) expand these metric functions. In the first step, the free functions should ideally cover the entire class of spacetimes without any redundancy. While this may seem trivial, a satisfactory answer with regards to all possible stationary and axisymmetric spacetimes is—to the best of our knowledge—not known. In Section IV.1.1, we review what is known in subclasses of stationary and axisymmetric spacetimes with additional symmetries. For the second step, a relation to physical quantities, such as multipole moments at asymptotic infinity, seems desirable. We briefly review such expansions in Section IV.1.2.
@@ IV.1.1 Subclasses of stationary and axisymmetric spacetimes with a redundancy-free metric ansatz
First, one may restrict to circular spacetimes [97, 98, 99]. Physically, circular spacetimes allow for a global foliation into two orthogonal 2-surfaces. This also implies a global symmetry under simultaneous inversion $\phi\leftrightarrow-\phi$ and $t\leftrightarrow-t$ of the two Killing coordinates (associated with $\eta^{\mu}$ and $\xi^{\mu}$). All vacuum solutions of GR are circular [^9] Mathematically, circularity can be understood as a property of the Ricci-tensor $R_{\mu\nu}$ in relation to the two Killing vectors $\eta^{\mu}$ and $\xi^{\mu}$, i.e., that $\eta^{\mu}R_{\mu}^{\,\,[\nu}\xi^{\kappa}\eta^{\lambda]}=0$ and $\xi^{\mu}R_{\mu}^{\,\,[\nu}\eta^{\kappa}\xi^{\lambda]}=0$ everywhere (while there exists a point in the spacetime for which $\eta^{[\mu}\xi^{\nu}\nabla^{\kappa}\eta^{\lambda]}=0$ and $\xi^{[\mu}\eta^{\nu}\nabla^{\kappa}\xi^{\lambda]}=0$), see e.g., [99, Sec. 7.1]. It is now obvious that any Ricci-flat spacetime, hence, any vacuum solution of GR is circular.. Irrespective of vaccum GR solutions, any circular spacetime can be written as
@@
where the choice of $\sigma(r,\theta)=r^{2}+a^{2}\cos(\theta)^{2}$ specifies Boyer–Lindquist coordinates $(t,r,\theta,\phi)$. The circular class requires five non-vanishing metric elements, described in terms of four independent functions [^10] Eq. (7) requires only four free functions but five non-vanishing metric elements. Hence, coordinate transformations can be used to, for instance, alter the relation between $dr^{2}$ and $d\theta^{2}$. This coordinate freedom can be used to reduce from five to four free functions, see, e.g., [97] for the well-known Papapetrou form. The latter choice of coordinates, however, does not contain Kerr in Boyer–Lindquist coordinates [96, App. B].. It is both minimal and exhaustive in the number of required free functions. Exhaustiveness of (7) relies on circularity and has been established in Refs. [97, 98, 99].
@@
Finally, in any of the above classes, one may or may not chose to break reflection symmetry about the equatorial plane, see e.g. [103] for a study within the Benenti–Francaviglia class. Of course, one may also further restrict to spherical symmetry, where stationarity implies staticity.
@@ IV.1.2 Expansion approaches to parametrized deviations
Multipole expansions of the Weyl class have been explored in Refs. [85, 104]. They have been extended to deviations from Kerr spacetime by applying the Newman-Janis algorithm [105] to the Weyl class [106, 104, 107, 108]. The resulting spacetimes are no longer within the Weyl class. In Ref. [107] it was shown that the resulting spacetimes do not necessarily admit for a rank-2 Killing tensor and thus need not remain within the Benenti–Francaviglia class.
Different Taylor expansions around asymptotic infinity within the Benenti–Francaviglia class have been explored in Refs. [109, 110, 111]. A continued fraction expansion around the horizon at $r=r_{0}$ of the circular class (first developed in spherical symmetry in Ref. [112]) has been set up by Konoplya, Rezzolla, and Zhidenko (KRZ) [113], which we review in Appendix B. To the best of our knowledge, the KRZ expansion covers all of the above expansions, while the reciprocal is not true. Hence, we focus on this continued-fraction expansion below.

### 05 Mar 2024 v2

@@ II.3 Resulting scheme and numerical implementation
Figure 3: Two schematic examples of $n=1$ lensing bands $\text{LB}^{n=1}(\mathfrak{g})$ (light-shaded regions), for which the respective spacetime $\mathfrak{g}$ remains allowed (left-hand panel) / can be excluded (right-hand panel), after a confident detection of a lensed emission region $\text{LB}^{n\geq 1}_{\text{VLBI}}$ (circular crescent; green regions contained in the lensing band; orange regions not contained in the lensing band).
From here on, we assume that a lensed emission region (i) has been observed and (ii) must lie, for any given spacetime $\mathfrak{g}$, within its $n=1$ lensing band, $\text{LB}^{n=1}(\mathfrak{g})$. Hence, if the observed lensed emission region $\text{LB}^{n\geq 1}_{\text{VLBI}}$ does not match, or “fit into,” the lensing-band $\text{LB}^{n=1}(\mathfrak{g})$, then we can rule out the respective underlying spacetime geometry $\mathfrak{g}$, see Fig. 3 for an schematic representation of this matching procedure. The level of confidence with which the underlying spacetime geometry can be ruled out depends on (i) the level of confidence in the detection of the lensed emission region, and (ii) the level of confidence in the theoretical assumption that it must occur within the lensing band.
We allow for an arbitrary centroid shift by an image-plane vector $(x_{0},y_{0})$ when performing the matching. The centroid shift amounts to the lack of a reference point. Potential observations of the inner shadow [30, 31, 32] or the base of the jet in M87* [82] could provide additional priors for the centroid shift. Examples of such reference points will be shown in Sec. IV.2, but we will not include them in the statistical analysis presented in this work. Such additional priors will only strengthen evidence against the underlying spacetime $\mathfrak{g}$.
In the following, we present an explicit algorithmic scheme with which we decide – given a lensed emission region $\text{LB}^{m\geq n}_{\text{VLBI}}$, as defined in Sec. II.2 – whether or not a given spacetime $\mathfrak{g}$ is excluded. The readers not interested in numerical technicalities may skip this section and proceed directly to the results (Sec. III).
@@ II.3.1 Free-floating bisection to obtain the lensing-band region
These conditions evaluate to TRUE whenever the respective image point is to the left of the respective lensing-band boundary $\partial^{\text{(outer)}}\text{LB}^{n}(\mathfrak{g})$ [or $\partial^{\text{(inner)}}\text{LB}^{n}(\mathfrak{g})$], where left and right are defined following the boundary counterclockwise. The only difference between determining the inner and outer boundary is a flip in the condition (or equivalently a flip in clockwise/counterclockwise direction). Other conventions may be chosen and are equivalent to the above.
@@
  • The second major advantage is the adaptive step size. The free-floating bisection initializes each subsequent bisection in reference to (at least two) previous points on the boundary. This makes it possible to dynamically adapt the step size and precision which, in turn, allows us to optimize efficiency while ensuring that the boundary is not lost.
We assume that the free-floating bisection algorithm can be adapted to any other (binary) boundary bisection condition in a 2D plane. Given the generality of this problem and given the above two advantages, the algorithm might be of more general use. We provide further details in App. A.
Given this free-floating bisection algorithm, the problem is reduced to standard numerical ray tracing (in curved spacetime $\mathfrak{g}$) at each bisection step. In order to keep the code-development minimal, we use an adapted version of Mathematica’s NDSolve routine for the ray tracing [^5] This is certainly not the most efficient ray tracing and we are inclined to make use of existing optimized ray tracing codes (such as the ones presented in Refs. [84, 16, 85, 86], for example) in future applications.. The ray tracing is performed in Boyer–Lindquist type coordinates. The translation from screen coordinates to ray tracing coordinates is detailed in App. C (see also Ref. [83]) and, in fact, only requires that the coordinates converge to oblate spheroidal coordinates at $r\rightarrow\infty$. The observers screen is then placed at $r=10^{5}\,M$ which approximates the $r\rightarrow\infty$ limit sufficiently well [^6] We have tested that the precision of the lensing-band boundary in the Kerr space time at high inclination (see Fig. 5 for some visual examples) remains at least at the subpercent level with the analytical results using the code described in Ref. [64].
@@ II.3.2 Minimizing the uncovered emission region
The normalization ensures that $0\leqslant\mathcal{A}(\mathfrak{g})\leqslant 1$ which is useful for numerical purposes. Once $\mathcal{A}$ is determined, we decide
@@ II.3.3 Bayesian inference
Given a confident measurement of lensed emission as discussed above, we estimate the parameters of the spacetime $\mathfrak{g}(M,A,p_{1},\ldots,p_{k})$, where $M$ and $A$ denote the asymptotic mass and spin, respectively, and $p_{i}$ denotes extra parameters (e.g., coupling constants or dimensionless “bumpy” parameters [88]), for which the $n=1$ lensing band of an observer at infinity, and at inclination $i$ can cover the considered ring. In other words, we estimate the set of parameters $\vec{\lambda}=\left\{i,M,A,p_{1},\ldots,p_{k}\right\}$ such that its $n=1$ lensing band covers the observed VLBI feature, as Fig. 3 illustrates.
@@
The emcee ensemble of walkers is initialized by sampling from the prior distributions. For all the cases presented in this work, we run until $\sim 10^{5}$ samples are obtained, and burn-in the initial $100$ samples. In general, parameter regions in which the lensing band is wide are easier to find within the posterior likelihood distribution. Recovering the theoretically expected symmetry of all results and correlations under the exchange of $A\leftrightarrow-A$ (as well as $a\leftrightarrow-a$) seems to be another good marker for convergence of the MCMC sampler. The presented results suggest that most (but not all) correlations are fully converged. Given the exploratory character of our work, we do not deem it justified to spend computation time on further convergence.
@@ III Benchmark: Projected constraints in the Kerr spacetime
We do, however, find an interesting correlation, albeit small, between mass and spin: the higher the absolute value of the spin, the larger the inferred mass, see the covariance of $A$ and $M$ in Fig.4. The underlying reason for this correlation is the shrinking of the $n=1$ lensing band with increasing spin, see left vs right columns in Fig. 5. With very different inference techniques, it has been previously found in Refs. [90, 40] that such correlation between mass and spin is more significant in the $n=1$ image than in the critical curve (i.e., in the $n=\infty$ image). It is reassuring that we recover this correlation.
As expected, given that the modeled VLBI feature is a circle, the results disfavor high inclinations. To be specific, the posterior for the projected constraint on inclination results $i\doteqdot{34^{\circ}}^{+28^{\circ}}_{-22^{\circ}}\;$. Indeed, the $n=1$ lensing band becomes very thin as the inclination approaches $i\sim 90^{\circ}$, as shown by the different rows in Fig. 5. We highlight that formally our analysis assumes, and thus all constraints rely on, a confident detection of lensed emission throughout the specified region $\text{LB}^{n\geq 1}_{\text{VLBI}}$, see, for instance, the darker and colored regions in Figs. 5 and 3 and the discussion in Subsec. II.2. In particular, the quantitative projected constraints (such as the one on the inclination) which arise from a thinning of the lensing band, implicitly rely on lensed emission being observed not just somewhere but everywhere within the observed region.
@@
Dedicated future studies to determine how confident VLBI observations can constrain the thickness of a lensed emission region are vital to solidify constraints such as the above ones. We also emphasize that the definition of lensing bands (defined with respect to the equatorial plane) is ill defined at exactly edge-on inclination $i=90^{\circ}$. Respectively, when deforming, e.g., the $n=1$ lensing band with increasing inclination across $i=90^{\circ}$, one encounters a discontinuity, see e.g., [45, Fig. 7]. The quantitative impact of this remains to be scrutinized in future studies and, hence, we are cautious to interpret our results at close to edge-on inclination.
@@ IV Application: Projected constraints on black hole non-uniqueness
  • efficiently explore degeneracies and correlations in large parameter spaces,
  • quantify the constraining power of a confident observation of a lensed emission region,
@@
As in the previous section, we present the results in the form of “projected constraints” on the underlying spacetime parameters. In Subsec. IV.1, we briefly review parametrized deviations in stationary and axisymmetric spacetimes beyond Kerr, and present a seven-parameter spacetime which we will use as a prototype test case beyond Kerr. In Subsec. IV.2, we then apply the lensing-band framework presented in Subsec. II and discuss the resulting projected constraints.
@@ IV.1 Parametrizing deviations to circular spacetimes
There are various ways in which black-hole candidates may differ from a black hole solution in GR [91]. In recent years, a “theory-agnostic” approach to testing GR has been heavily pursued. This approach involves parametrizing deviations from GR solutions (such as the metric [88] or gravitational potentials [8]) or observables (such as the gravitational waveform [92] or quasinormal modes frequencies [93]). As a result, a particular theory is neither employed nor tested, but the hope is that the parameters involved, in such an agnostic way, can then be mapped to physical constraints.
The exploration of deviations at the level of the metric requires to (i) parametrize a given class of spacetimes in terms of free metric functions, and then (ii) expand these metric functions. In the first step, the free functions should ideally cover the entire class of spacetimes without any redundancy. While this may seem trivial, a satisfactory answer with regards to all possible stationary and axisymmetric spacetimes is—to the best of our knowledge—not known. In Subsec. IV.1.1, we review what is known in subclasses of stationary and axisymmetric spacetimes with additional symmetries. For the second step, a relation to physical quantities, such as multipole moments at asymptotic infinity, seems desirable. We briefly review such expansions in Subsec. IV.1.2.
@@ IV.1.1 Subclasses of stationary and axisymmetric spacetimes with a redundancy-free metric ansatz
First, one may restrict to circular spacetimes [100, 101, 102]. Physically, circular spacetimes allow for a global foliation into two orthogonal 2-surfaces. This also implies a global symmetry under simultaneous inversion $\phi\leftrightarrow-\phi$ and $t\leftrightarrow-t$ of the two Killing coordinates (associated with $\eta^{\mu}$ and $\xi^{\mu}$). All vacuum solutions of GR are circular [^9] Mathematically, circularity can be understood as a property of the Ricci-tensor $R_{\mu\nu}$ in relation to the two Killing vectors $\eta^{\mu}$ and $\xi^{\mu}$, i.e., that $\eta^{\mu}R_{\mu}^{\,\,[\nu}\xi^{\kappa}\eta^{\lambda]}=0$ and $\xi^{\mu}R_{\mu}^{\,\,[\nu}\eta^{\kappa}\xi^{\lambda]}=0$ everywhere (while there exists a point in the spacetime for which $\eta^{[\mu}\xi^{\nu}\nabla^{\kappa}\eta^{\lambda]}=0$ and $\xi^{[\mu}\eta^{\nu}\nabla^{\kappa}\xi^{\lambda]}=0$), see e.g., Ref. [102, Sec. 7.1]. It is now obvious that any Ricci-flat spacetime, hence, any vacuum solution of GR is circular.. Irrespective of vacuum GR solutions, any circular spacetime can be written as
@@
where the choice of $\sigma(r,\theta)=r^{2}+a^{2}\cos(\theta)^{2}$ specifies Boyer–Lindquist coordinates $(t,r,\theta,\phi)$. The circular class requires five non-vanishing metric elements, described in terms of four independent functions [^10] Equation. (7) requires only four free functions but five non-vanishing metric elements. Hence, coordinate transformations can be used to, for instance, alter the relation between $dr^{2}$ and $d\theta^{2}$. This coordinate freedom can be used to reduce from five to four free functions, see, e.g., Ref. [100] for the well-known Papapetrou form. The latter choice of coordinates, however, does not contain Kerr in Boyer–Lindquist coordinates [99, App. B].. It is both minimal and exhaustive in the number of required free functions. Exhaustiveness of Eq. 7 relies on circularity and has been established in Refs. [100, 101, 102].
@@
Finally, in any of the above classes, one may or may not chose to break reflection symmetry about the equatorial plane, see e.g., Ref. [106] for a study within the Benenti–Francaviglia class. Of course, one may also further restrict to spherical symmetry, where stationarity implies staticity.
@@ IV.1.2 Expansion approaches to parametrized deviations
Multipole expansions of the Weyl class have been explored in Refs. [88, 107]. They have been extended to deviations from Kerr spacetime by applying the Newman–Janis algorithm [108] to the Weyl class [109, 107, 110, 111]. The resulting spacetimes are no longer within the Weyl class. In Ref. [110] it was shown that the resulting spacetimes do not necessarily admit for a rank-2 Killing tensor and thus need not remain within the Benenti–Francaviglia class.
Different Taylor expansions around asymptotic infinity within the Benenti–Francaviglia class have been explored in Refs. [112, 113, 114]. A continued fraction expansion around the horizon at $r=r_{0}$ of the circular class (first developed in spherical symmetry in Ref. [115]) has been set up by Konoplya, Rezzolla, and Zhidenko (KRZ) [116], which we briefly review in App B. To the best of our knowledge, the KRZ expansion covers all of the above expansions, while the reciprocal is not true. Hence, we focus on this continued-fraction expansion below.